# A Prompt-Engineered Large Language Model, Deep Learning Workflow for Materials Classification

### v1

@@ Abstract
With the advent of ChatGPT, large language models (LLMs) have demonstrated considerable progress across a wide array of domains. Owing to the extensive number of parameters and training data in LLMs, these models inherently encompass an expansive and comprehensive materials knowledge database, far exceeding the capabilities of individual researcher. Nonetheless, devising methods to harness the knowledge embedded within LLMs for the design and discovery of novel materials remains a formidable challenge. In this study, we introduce a general approach for addressing materials classification problems, which incorporates LLMs, prompt engineering, and deep learning algorithms. Utilizing a dataset of metallic glasses as a case study, our methodology achieved an improvement of up to 463% in prediction accuracy compared to conventional classification models. These findings underscore the potential of leveraging textual knowledge generated by LLMs for materials especially with sparse datasets, thereby promoting innovation in materials discovery and design.
@@
The concept of “AI for Materials” refers to the utilization of artificial intelligence (AI) techniques, such as machine learning (ML), deep learning (DL), and the increasingly prevalent large language models (LLMs), to design novel materials or investigate the composition-structure-property relationships in different materials [1, 2, 3, 4]. Significant strides have been achieved in fields such as biomaterials [5, 6] and organic materials [7, 8], which can be attributed to the relatively uniform representation of organic molecules [9] and the availability of numerous high-quality datasets [10]. However, for inorganic materials, particularly those with large compositional space (element types 4), the application of AI presents a more complex challenge. This complexity arises due to factors such as the scarcity of experimental data [4], diverse properties of inorganic materials, complicated crystal structures, potential existence of intermediate phases (e.g. intermetallics in alloys), and defects including vacancies, dislocations, and grain boundaries [11, 12]. Contrasting with organic molecules that can be modeled with neural networks, multi-component inorganic materials often involve the amalgamation of various phases and defects or even the formation of new phases, presenting an ongoing challenge in the structural representation [13, 14]. Although large-scale databases such as the Materials Project [15], ICSD [16], and AFLOW [17] exist, data pertinent to specific tasks such as the relationships between composition and mechanical properties in alloys and structural stability of two-dimensional materials are often sparse and dispersed across different datasets, complicating the assembly of substantial databases for training purposes. Furthermore, input features can vary across datasets, sometimes necessitating manual feature construction based on domain knowledge [18]. Collectively, these issues hinder the application of AI in the realm of inorganic materials, subsequently impeding the discovery and design of new materials.
@@
The advancement of LLMs, particularly with the advent of ChatGPT, has sparked a surge of interest in distilling knowledge from these models through prompt engineering [19, 20, 21]. Notable examples include constructing scientific question answering knowledge bases using LLMs [22], transferring knowledge from large to small models via in-context learning [23], and training materials-specific LLMs [24, 25]. Given the expanding capabilities of LLMs and the growing volume of trained data, it is conceivable to consider these models as encyclopedia resources for materials science, enabling the extraction of text-based knowledge [26]. As shown in the upper part of Fig. 1, this approach addresses some challenges associated with data collection and feature extraction in conventional processes. By representing all content in textual format, the generated data becomes universally applicable, offering a versatile solution for materials science research.
@@
Therefore, we have established a universal workflow for material text feature-label classification using textual data generated by LLMs, as illustrated in the lower part of Fig. 1. The workflow is structured as follows: (i) define the material classification problem to be addressed; (ii) design prompts via prompt engineering to distill knowledge from LLMs and store it as textual data; (iii) fine-tune a bidirectional encoder representations transformers (BERT) [27] model (commonly employed in natural language processing) to train on the textual data-label pairs; (iv) apply the model to explore new materials or study composition-structure-property relationships. In an example problem involving the classification of 5,577 metallic glasses (MGs) with experimental datasets labeling the MG forming categories (bulk MG, ribbon, or non-ribbon), a BERT model trained with optimized prompts and position embedding methods achieved up to a 48% increase in classification accuracy compared to traditional ML models. The classification models obtained from our workflow achieved an overall accuracy of 97.7%. For the classification of bulk MGs (BMGs) with the smallest sample number (11% of the entire dataset), the accuracy improved by up to 463% compared to traditional ML models. These results underscore the superiority of our workflow in addressing material classification problems and this workflow can be extended to various material applications. Additionally, tremendous potential of “AI for Materials” is revealed, grounded in natural language processing and prompt engineering.
Text-based materials
classification workflow
@@
Our material classification workflow comprises the four-tier structure depicted within the blue box in Fig. 1, and is readily adaptable to various material classification problems. Initially, researchers are required to undertake fundamental data collection and delineate the prediction objectives and categories, such as determining whether a material exhibits stability, corrosion resistance, or favorable optoelectronic transmission properties.
@@
Secondly, the inputs for ML models in conventional workflows typically comprise material properties obtained through experiments or theoretical calculations, demanding substantial expertise as well as experimental or simulation efforts. Different researchers may select varying property features, leading to multiple datasets for the same problem [28]. However, training numerical-based ML models on the combination of different datasets often fails due to inconsistencies in input feature vectors. Furthermore, the number of datasets for specific niche tasks in materials science research is quite small ( several hundreds), negligible compared to the millions of samples in classical computer science fields like ImageNet dataset [29]. For example, in the field of alloys, the dataset for fracture and impact toughness of high-entropy alloys contains only 154 data points [30], the fatigue database of complex metallic alloys consists of merely 272 data points [31], and the dataset of mechanical properties of high-entropy alloys has 370 data points [32]. These limitations hinder the application of ML in alloys. In our workflow, we employ textual inputs generated by LLMs, offering the advantage of eliminating the need for manual feature extraction. Moreover, LLMs can output content relevant to specific problems in materials science with customized prompts. As LLMs possess considerably more parameters than traditional ML models, they can generate a richer and more diversified knowledge base.
@@
Furthermore, within our workflow, the classification BERT model employs the attention mechanism for self-instruct learning, enabling automatic feature extraction. Different pre-trained models can be utilized for corresponding downstream tasks. Models based on the attention mechanism are not limited to text input but also applicable to areas like image recognition [33] and molecular structure identification [34]. These features allow our workflow to seamlessly support the combination of different datasets and multimodal expansion, addressing the issue of insufficient datasets for specific material problems often encountered in traditional methods.
@@
Finally, the material classification models trained through our workflow can be utilized in numerous downstream tasks, such as discovering new materials across different categories, conducting interpretability analysis to extract composition-structure-property relationships of materials, and even fine-tuning models to achieve excellent results in other classification tasks. Our workflow represents a novel and general approach for material sciences by constructing large-scale databases from prompt-engineered LLMs and training general DL models to expedite materials discovery. We showcase the efficiency and accuracy of the workflow by using the classification problem of MGs as an example.
Application: metallic glasses classification
@@
MGs are amorphous materials typically composed of a combination of metal and other elements [35]. They are renowned for superior engineering performance, including high strength, high thermal stability, and excellent corrosion resistance [36]. Despite ongoing efforts to discover MGs with enhanced comprehensive performance, their formability remains a complex and challenging problem in materials science. Predicting the glass formability is also difficult for atomistic simulations, due to the large size and long timescale (unattainable by first-principles calculations) and the absence of accurate interatomic potentials for molecular dynamics simulations. Moreover, MGs have an extensive compositional space, often comprising three or more elements, rendering the experimental search for new applicable MGs a time-consuming process [37, 38].
@@
Various ML methods [39] have been employed to classify MGs; however, these methods often rely on manually calculated or carefully designed material features. The inconsistency of designed features and datasets across different ML models necessitates researchers to recalculate those features each time when utilizing other models. In the worst-case scenario, certain features rely on experimental calculations, rendering a MG dataset applicable to one model but not another. In this regard, we leveraged the experimental dataset of MGs by [40] and applied it to the proposed workflow here. The dataset compiled by [40] consists of 8,415 alloy compositions with their corresponding MG forming categories. The alloy compositions were classified into three categories based on the critical cooling rate (): BMG (), ribbon (), and non-ribbon (NR, ). The glass formability is then BMG ribbon NR. From this dataset, 5,577 alloy compositions with only a single glass-forming category were selected as input dataset, with 80% of them being train set while the rest being test set. As shown in Fig. 2a, a customized version of the general workflow was developed for the given input data. The blue arrows indicate the training process, which involves expanding the alloy compositions into textual data using an LLM. Subsequently, a DL model is trained on the textual data from the training set for MG classification. Finally, the analysis of efficiency and accuracy and model interpretation are performed using textual data from the test set. These three main steps are illustrated in Fig. 2b. The green arrows indicate the application stages of the model. For an unexplored alloy composition, the workflow can determine the MG category (BMG, ribbon, or NR) within seconds.
@@
Next, we trained a MGs BERT (MgBERT) classification model that combines self-instruct learning and supervised classification. DL models for natural language processing typically require a vast number of fitting parameters, necessitating the selection of a suitable pre-trained model as the base model. MgBERT is fine-tuned based on MatSciBERT (a pre-trained model based on materials science texts) [42]. Figs. 4a, b display a comparison of different pre-trained BERT models, with MatSciBERT achieving the best results in classification accuracy and model loss on the test set. However, MatSciBERT can only accept inputs of up to 512 tokens, while our text data has a maximum length of 859 when converted to tokens. Consequently, we designed a dynamically resizable embedding layer called “MgBERT Embedding” to meet the requirements of input scalability. This embedding layer can assign weights based on layers and proportions according to the preset length of the new token, building upon the embedding weights trained by MatSciBERT. This approach generates new position embedding that can support longer text inputs and circumvent time-consuming weight initialization training. This new position embedding layer (in Fig. 4c) is utilized for tokenizing the input text of the model. From the comparison of results in Figs. 4g and h, our MgBERT embedding Method 3 (Fig. 4f) exhibits higher accuracy and lower train set loss, compared to the embedding layer method in Fig. 4d, which uses newly initialized weights, and that in Fig. 4e, which only employs pre-trained weights to cover the first 512 embedded positions. For different pre-trained models trained on the same text data (Figs. 4i, j), our MgBERT has the best performance in training set loss, training set classification accuracy, and test set classification accuracy. Notably, for alloy data in the test set that our model has never encountered before, our model achieved an accuracy of 88.5%, an improvement of 9% compared to Longformer, and an improvement of 4% compared to the best performing MatSciBERT. The detailed classification results can be observed in the confusion matrix in Supplementary Fig. S6. These findings showcase the success of our improvement in embedding structure, as it simultaneously supports longer input text and enhances accuracy. To examine the effectiveness of the entire workflow, we trained a baseline model based on logistic regression (LR) using the training set, as well as a support vector machine (SVM) model and a gradient boosting decision tree (GBDT) model. Then, we conducted tests on the test set and compared the classification performance of different methods (Fig. 5). For the default three-class problem, our workflow achieved a maximum accuracy increase of 48% across the entire dataset and a peak enhancement of 32% on the test set. Even more astonishing is our performance with respect to identifying BMGs, as shown in Fig. 5b. For this task, our accuracy soared by up to 463% on the entire dataset and by as much as 307% on the test set. Considering that BMGs constitute only 11% of our dataset, these results highlight the remarkable capability of our text-based material classification workflow in accurately recognizing patterns in data-scarce scenarios, surpassing traditional ML models.
@@
DL models are often considered to lack algorithmic transparency [43]. Language models such as BERT typically have hundreds of millions of parameters [27], making it challenging to provide detailed explanations of the model at the algorithmic level. Here, we utilized the local interpretable model agnostic interpretations (LIME) [44] method to analyze from an input-output perspective. We selected text data of alloy compositions CuZrGa (classified as BMG), AgAlLa (classified as ribbon), and AlMnSi (classified as NR) as three examples. Fig. 6a and Supplementary Fig. S7 illustrate that alloy composition is the primary contributing factor to these classification results, consistent with the classical Inoue rule [45] and high-throughput experiment results [46]. According to Inoue’s rule, components of an alloy introduce varying atomic size differences, altering the internal dense random packing structure. Moreover, the model has extracted some elemental property information. For example, “Melting” refers to the melting point, which is one of the critical criteria for determining the formation of MGs [47]. Additionally, “745” is the ionization energy of copper, which could affect the interatomic forces within the alloy and thereby influence MG formation. Furthermore, “2477” represents the boiling point of gallium, and the polymer-like thermoplastic behavior observed in MGs is thought to be related to boiling points [48], also connected to the formation of functional MGs. These findings indicate that our text classification model can perform self-instruct learning and automatically extract features from a given text. Building on this foundation, in contrast to traditional ML models that rely on manually designed features as input, our text-based workflow enables models to accept raw text as input. This flexibility allows for rapid transfer and recycling of pre-trained model parameters between different materials classification tasks. We also visualized the attention scores within the classification model and the model extracts features based on three distinct patterns (Fig. 6b). This process ensures that after 12 layers of training, the output [CLS] (classify) token possesses sufficient information for classification. Since we use the value of the [CLS] token from the final layer to make judgments in the classification layer, Fig. 6c visualizes the emphasis placed on different sections of the text. The results suggest that the composition and alloy layers have garnered significant attention. A possible explanation is that key tokens within these two layers have captured essential features from the input text. Consequently, the [CLS] token can achieve accurate classification results by focusing on those specific token positions. These findings demonstrate that the classification MgBERT model, generated by our general workflow, has the abilities of dynamic feature extraction, modeling inter-word relationships, hierarchical representation learning, and bidirectional context understanding, which may be potential indications for good classification performance.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated rapid progress across a wide array of domains. Owing to the very large number of parameters and training data in LLMs, these models inherently encompass an expansive and comprehensive materials knowledge database, far exceeding the capabilities of individual researcher. Nonetheless, devising methods to harness the knowledge embedded within LLMs for the design and discovery of novel materials remains a formidable challenge. We introduce a general approach for addressing materials classification problems, which incorporates LLMs, prompt engineering, and deep learning. Utilizing a dataset of metallic glasses as a case study, our methodology achieved an improvement of up to 463% in prediction accuracy compared to conventional classification models. These findings underscore the potential of leveraging textual knowledge generated by LLMs for materials especially in the common situation where datasets are sparse, thereby promoting innovation in materials discovery and design.
@@
The concept of “AI for Materials” refers to the utilization of artificial intelligence (AI) techniques, such as machine learning (ML), deep learning (DL), and the increasingly prevalent large language models (LLMs), to design novel materials and to investigate composition-structure-property relationships [1, 2, 3, 4]. Significant strides have been achieved in fields such as biomaterials [5, 6] and organic materials [7, 8] using such approaches, which can be attributed to the relatively simple representation of the structures of organic molecules [9] and the availability of numerous high-quality datasets [10]. However, for inorganic materials, particularly those with large compositional space (element types 4), the application of AI presents a more complex challenge. This complexity arises due to factors such as the scarcity of experimental data [4], the diversity of material properties of interest, the complexity of crystal structures, the existence of multiple alloy phases, and the complexity/properties of defects [11, 12] (e.g., vacancies, dislocations, grain boundaries). In contrast with organic molecules that can be naturally modeled using neural networks, multi-component inorganic materials often involve the amalgamation of multiple phases and defects or even the formation of new phases, presenting an ongoing challenge for structural representation [13, 14]. Although large-scale databases such as the Materials Project [15], ICSD [16], and AFLOW [17] exist, data pertinent to specific tasks such as the relationships between composition and mechanical properties in alloys and structural stability of two-dimensional materials are often sparse and dispersed across different datasets, complicating the assembly of large scale training sets. Input features often vary between datasets, necessitating manual feature construction based on domain knowledge [18]. Collectively, these issues have hindered AI applications in inorganic materials, subsequently impeding the discovery and design of new materials.
@@
The advancement of LLMs, particularly ChatGPT, sparked a surge of interest in distilling knowledge through prompt engineering [19, 20, 21]. Notable examples include constructing scientific question-answering knowledge bases using LLMs [22], transferring knowledge from large to small models via in-context learning [23], and training materials-specific LLMs [24, 25]. Given the expanding capabilities of LLMs and the growing volume of training data, it is conceivable to view these models as encyclopedic resources for materials science, enabling the extraction of text-based knowledge [26]. As shown in the upper part of Fig. 1, this approach addresses some challenges associated with data collection and feature extraction in classical ML processes. By representing content in textual format, the generated data becomes universally applicable, offering a versatile tool for materials science research.
@@
Here, we establish a universal workflow for material text feature-label classification using textual data generated by LLMs, as illustrated in the lower part of Fig. 1. The workflow is structured as follows: (i) define the material classification problem to be addressed; (ii) design prompts (via prompt engineering) to distill knowledge from LLMs and store the information as textual data; (iii) fine-tune a bidirectional encoder representations from transformers (BERT) [27] model (commonly employed in natural language processing) to train on the stored textual data-label pairs; (iv) apply the model to explore new materials or study composition-structure-property relationships. In an example problem involving the classification of 5,577 metallic glasses (MGs) with experimental datasets labeling MG-formability categories (bulk MG, ribbon, or non-ribbon – i.e., glass formation under different cooling rates), a BERT model trained with optimized prompts and position embedding methods achieved up to a 48% increase in classification accuracy compared to traditional ML models. The classification models obtained from the workflow presented here achieved an overall accuracy of 97.7%. For the classification of bulk MGs (BMGs - those formed at the slowest cooling rates) from the smallest training dataset (11% of the entire dataset), the accuracy improved by up to 463% compared to traditional ML models. These results underscore the superiority of the proposed workflow in addressing material classification problems. Since this workflow is extendable to a wide range of material applications, this approach demonstrates the tremendous potential of “AI for Materials” grounded in natural language processing and prompt engineering for materials engineering.
Text-based materials classification workflow
@@
Our material classification workflow has four tiers depicted in the blue box in Fig. 1, and is readily adaptable to various material classification problems. Initially, researchers undertake fundamental data collection and delineate the prediction objectives and categories, such as determining whether a material exhibits stability, corrosion resistance, or favorable optoelectronic transmission properties.
@@
The inputs to conventional ML model workflows typically comprise material properties (obtained through experiments or theoretical calculations), demanding substantial expertise and effort to obtain experimental or simulation data. Different researchers may focus on different input features, leading to multiple datasets for the same problem [28]. Training numerical-based ML models on a combination of different datasets often fails due to inconsistencies in input feature vectors. The size of datasets for specific niche tasks in materials science research tends to be relatively small ( several hundred entries); this is negligible compared to the millions of samples in classical computer science fields like the ImageNet dataset [29]. For example, in the field of alloys, a typical set of data for the fracture and impact toughness of high-entropy alloys contains only 154 entries [30], the fatigue database of complex metallic alloys consists of 272 entries [31], while other mechanical properties of high-entropy alloys tabulations contain 370 entries [32]. These limitations hinder the application of ML in alloys. In our workflow, we employ textual inputs generated by LLMs, offering the advantage of eliminating the need for manual feature extraction. The second tier in our workflow is the application of LLMs to output content relevant to specific problems in materials science with customized prompts. As LLMs possess considerably more parameters than traditional ML models, they generate a richer and more diversified knowledge base.
@@
In the third tier, a BERT model, employing an attention mechanism for Self-Instruct learning approach for automatic feature extraction, provides classification. Different pre-trained models are utilized for different downstream tasks. Models based on the attention mechanism are not limited to text input but may also be applied to areas like image recognition [33] and molecular structure identification [34]. These features allow our workflow to seamlessly support a combination of different datasets and multimodal expansion, addressing the issue of insufficient datasets for specific material problems often encountered in traditional methods.
@@
Finally, the material classification models trained through our workflow can be applied to numerous downstream tasks, such as discovering new materials for different technologies, conducting interpretability analysis to extract composition-structure-property relationships, and even fine-tuning models to achieve even better results in other classification tasks. Our workflow represents a novel, general approach for material sciences by constructing large-scale databases from prompt-engineered LLMs and training general DL models to expedite materials discovery. We showcase the efficiency and accuracy of the workflow by applying it to the MG classification problem described above.
Application: metallic glass classification
@@
MGs are amorphous structures typically containing several metals and non-metallic elements [35]. The family of MGs is renowned for superior engineering performance, including high strength, high thermal stability, and excellent corrosion resistance [36]. Despite ongoing efforts to discover MGs with enhanced comprehensive performance, their formability remains a complex and challenging problem in materials science. Predicting glass formability (thermal quench rates required for glass formation) is also difficult for atomistic simulations due to large sample size and, especially, the long timescale (inaccessible to first-principles calculations) and the absence of accurate interatomic potentials for molecular dynamics simulations. Moreover, the compositional space of MGs is of high dimension, often comprising more than three elements, rendering the experimental search for new applicable MGs a time-consuming process [37, 38].
@@
Various ML methods [39] have been employed to classify MGs, usually based upon manually calculated or carefully designed input material features. The inconsistency of designed features and datasets across different ML models often necessitates recalculating those features when utilizing other models. In the worst case, experiments must be rerun to provide input to different models. In this regard, we leveraged existing MG experimental data [40] and applied it to the proposed workflow here. The dataset compiled by [40] consists of 8,415 alloy compositions classified according to MG formability. The alloy compositions were classified into three categories based on the critical cooling rate (), BMG (), ribbon (), and non-ribbon (NR, ); i.e., the glass formability is then BMG ribbon NR. From this dataset, we focus on 5,577 alloy compositions, of which 80% were used in the training set - the remainder forming the test set. As shown in Fig. 2a, a customized version of the general workflow was developed for this input data type. The blue arrows indicate the training process, which involves expanding the information on each alloy (textual data - e.g., physical properties) using an LLM. Subsequently, a DL model is trained on the textual data from the LLM-expanded training set for MG classification. Finally, the efficiency and accuracy are analyzed using the textual data from the test set. We then query the model to interpret the results. The green arrows in Fig. 2(a) indicate the application stages of the model; the three main steps (below) are illustrated in Fig. 2(b). The workflow can determine the MG category (BMG, ribbon, or NR) within seconds for an unexplored alloy composition.
@@
Next, we trained an MG BERT (MgBERT) classification model combining Self-Instruct learning and supervised classification. DL models for natural language processing typically require a vast number of fitting parameters, necessitating the selection of a suitable pre-trained model as a base. MgBERT is fine-tuned based on the MatSciBERT (a pre-trained model based on materials science texts) [42]. Figs. 4(a, b) compare different pre-trained BERT models, with that based upon MatSciBERT achieving the best classification accuracy and model loss on the test set. While MatSciBERT only accepts inputs of up to 512 tokens, our text data has a maximum length of 859 when converted to tokens. Consequently, we designed a dynamically resizable embedding layer called “MgBERT Embedding” to meet the input scale requirements.
This embedding layer assigns weights based on layers and proportions according to the preset length of the new token, building upon the embedding weights trained by MatSciBERT. This approach generates new position embeddings that support longer text inputs and circumvent time-consuming weight initialization training. This position embedding layer (in Fig. 4(c)) is employed to tokenize the input text of the model. Comparison of the results in Figs. 4(g) and (h) show that the MgBERT embedding Method 3 (Fig. 4(f)) exhibits higher accuracy and lower training set loss, compared to the embedding layer method in Fig. 4(d) (using newly initialized weights), and directly using pre-trained weights (covering the first 512 embedded positions - see Fig. 4(e)). For different pre-trained models based on the same text data (Figs. 4(i, j)), our MgBERT has the best performance in terms of the training set loss, training set classification accuracy, and test set classification accuracy. Notably, for alloy data in the test set (not used in the training), our model achieved an accuracy of 88.5% (an improvement of 9% compared to Longformer and an improvement of 4% compared to the best performing MatSciBERT). The detailed classification results are shown through a confusion matrix in Supplementary Fig. S6. These findings showcase the success of our improvement in embedding structure, simultaneously supporting longer input text and enhanced accuracy.
To examine the effectiveness of the entire workflow, we trained a baseline model based on logistic regression (LR) using the training set, as well as a support vector machine (SVM) model and a gradient boosting decision tree (GBDT) model. We tested the results (on the test set) and compared the classification performance of different methods (Fig. 5). For the default three-category MG classification problem, our workflow achieved a maximum accuracy increase of 48% across the entire dataset and a peak enhancement of 32% on the test set (as compared with the baseline). The accuracy of the workflow in identifying BMGs, as shown in Fig. 5b, increased by 463% on the entire dataset and by as much as 307% on the test set as compared with the baseline (traditional ML) classification model. Considering that BMGs constitute only 11% of our dataset, these results highlight the capability of our text-based material classification workflow in accurately recognizing patterns in data-scarce scenarios, far surpassing traditional ML models.
DL models often lack algorithmic transparency [43]. Language models such as BERT typically have hundreds of millions of parameters [27], making it challenging to explain the model results (at the algorithmic level). We applied the local interpretable model-agnostic interpretations (LIME) [44] method to analyze the input-output relationship. We selected text data of alloy compositions Cu55Zr42.5Ga2.5 (BMG), Ag20Al25La55 (ribbon), and Al40Mn25Si35 (NR) for this analysis. Fig. 6(a) and Supplementary Fig. S7 show to what degree the words in the input data from the LLM (marked LLM Output in Fig. 3) are most important for the input/output correlation. Fig. 6(a) shows that the maximum contribution to this correlation is the composition Cu55Zr42, with a value of +0.88. On the other hand, for the Ag20Al25La55 case, the alloy contribution to the correlation is much weaker. Since the system “knows” the atomic sizes and that those atomic sizes are much different in the BMG CuZr alloy cases (compared with the ribbon AgAlLa case), this may suggest that atomic size difference is important, consistent with the classical Inoue rule [45] for glass formability and high-throughput experiment results [46] (i.e., easy glass formation for more than three atom types with atomic size ratios above 12%). The model also shows that some properties of individual elements are important. For example, “Melting” refers to elemental melting points, which is also used as one of the parameters commonly employed as a predictor for glass formation [47]. The word “745” in the CuZr alloy in Fig. 6(a) refers to the ionization energy of copper - this may influence the tendency for charge transfer between Cu and other alloying elements. The word “2477” represents the boiling point of gallium; elemental boiling points have been suggested as related to some MG behavior [48].
@@
These findings suggest that our text classification approach can be used to automatically extract features from the textual input. Building on this foundation, in contrast to traditional ML models that rely upon manually designed features as input, our text-based workflow enables models to accept any raw text as input. This flexibility allows for rapidly transferring and recycling pre-trained model parameters between different materials classification tasks. We also visualized the attention scores within the classification model, and the model extracts features based on three distinct patterns (Fig. 6(b)). An attention score tells the importance of the contribution of each “neuron” strength in this 12-layer deep neural net. This process ensures that after 12 layers of training, the output [CLS] (classify) token possesses sufficient information for classification. Since we use the value of the [CLS] token from the final layer to make judgments in the classification layer, Fig. 6(c) visualizes the importance of different sections of the input text. The results suggest that the composition and alloy layers have garnered significant attention (compared with the element layer - element information is implicitly contained in the other two layers). A possible explanation is that key tokens within these two layers have captured the essential features from the input text. Consequently, the [CLS] token can achieve accurate classification results by focusing on those specific token positions. These findings demonstrate that the classification MgBERT model, generated by our general workflow, has the abilities of dynamic feature extraction, modeling inter-word relationships, hierarchical representation learning, and bidirectional context understanding, which are indicators of good classification performance.